In six attempts it took a locksmith 9, 14, 7, 8 11, 5 seconds to open a certain kind of lock. Construct a 95% confidence interval for the average it takes the locksmith to open this kind of lock.

1
Expert's answer
2021-01-18T19:29:46-0500

"\bar{x} = \frac{9 + 14 +7+8+11+5}{6} \\

= 9"

Standard deviation

"SD = 3.1623 \\

n = 6"

Degree of freedom

"df = n - 1 \\

= 6 - 1 \\

= 5"

we look into t-distribution table with df and with 95% confidence level

Critical value t = 2.571

Confidence interval formula:

"= \bar{x} ± \frac{t \times SD}{\sqrt{n}} \\

= 9 ± \frac{2.571 \times 3.1623}{\sqrt{6}} \\

= 9 ± 3.32 \\"

5.68 < μ < 12.32